OpenStudy (anonymous):

Which ordered pair is a solution of the equation 2x − y = 9? (-4,1) (-2,5) (5,1) (6,-3) @kelliegirl33

OpenStudy (anonymous):

just a matter of subbing...your points are ALWAYS set up as (x,y) 2x - y = 9 lets try the first answer choice...(-4,1)...x = -4 and y = 1 2(-4) - 1 = 9 -8 - 1 = 9 -9 = 9 (incorrect) so this is not the one you want to try the 2nd answer choice ?

OpenStudy (anonymous):

(-2,5) is the answer?

OpenStudy (anonymous):

it is also easy to check.... (-2,5) 2x - y = 9 2(-2) - 5 = 9 -4 - 5 = 9 -9 = 9 (incorrect) no...(-2,5) is not the answer

OpenStudy (anonymous):

i think it is prob C?

OpenStudy (anonymous):

lets check.. 2x - y = 9.....(5,1) 2(5) - 1 = 9 10 - 1 = 9 9 = 9 (correct) BINGO...it is C

OpenStudy (anonymous):

Thank you so much for all your help!

OpenStudy (anonymous):

anytime :)